OpenStudy (theislander):

Find the equation in Point-Slope form of the line satisfying the conditions: Through (3,-5) and (-2,1)

OpenStudy (anonymous):

first you find the slope \[\frac{ 1-(-5)}{ -2-3 }\] \[-\frac{ 6 }{ 5 }\]

OpenStudy (anonymous):

point slope is\[y-y _{1}=m(x-x _{1})\]\[y-1=-\frac{ 6 }{ 5 }(x-(-2))\]

OpenStudy (anonymous):

distribute the -6/5, then get y by itself by adding one to both sides.